[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jul 03, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Jul 3 10:17:38 UTC 2016


HSU Shift Log for Sun Jul 03 2016 06:17:32 GMT-0400 (EDT)

Shift: A

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS/AWIPS2 issues: METEOSAT pseudo-color satellite products have not  
been ingesting since about 0430Z 02 July.

Other Techincal issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160703/98ae7683/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list